graphing equations with two variables. y= -1/2x +3
Answer by checkley71(8403)

WHAT YOU HAVE IS A STANDARD LINE FORMULA Y=mX+b WHERE m=THE SLOPE (-1/2) & b=THE Y INTERCEPT (0,3).
THE SLOPE IS FOUND BY THE FOLLOWING FRACTION (Y2-Y1)/X2-X1).
TO FIND A SECOND POINT USING THIS EAUATION WE SET Y=0 & SOLVE FOR X THUS
0=-1/2X+3
1/2X=3
X=3*2
X=6 OR (6,0)
NOW PLOTING THESE TWO POINTS (0,3) &(6,0) & DRAWING A LINE THROUGH THEM YOU HAVE A SLOPE OF -1/2.
HAPPY GRAFTING :)